# Heart Failure Prediction using Modal Decomposition and Masked Autoencoders for Scarce Echocardiography Databases

### v1

@@ 5 Conclusions
Heart diseases are the main cause of human mortality in the world. The early identification of heart diseases is thus a task of great importance. In this work, a real-time heart failure prediction system for echocardiography video sequences has been presented to specifically address this challenging and demanding task, which has not been addressed in the related literature to the best of the authors’ knowledge. The two major contributions of this paper are the creation of a large annotated cardiac database from echocardiography sequences with heterogeneous acquisition features; and a deep neural network, based on a ViT, trained using a self-supervised learning approach which effectively improves the performance with scarce datasets. These two contributions, acting synergistically, successfully address the usual challenge of having a scarce number of samples in the medicine field, reducing the dependence on very large training datasets. More precisely, the elaboration of varied databases with high quality enough is very difficult, implying high costs and very hard specialized work from experts in heart diseases. In addition, the general trend in the related literature about heart disease recognition and heart failure prediction in echocardiography images barely explores self-supervised learning, even less the Masked Autoencoder (MAE), to deal with this typical situation of scarce databases. This implies the necessity to adopt techniques specifically designed to address this problem. In the creation process of the database, the HODMD algorithm has been employed as a feature extractor and as a data augmentation technique. This algorithm has demonstrated to improve the heart failure time prediction performance, as also incorporating temporal information from the HODMD modes, useful to better characterize the different heart states and so to more accurately estimate the times of heart failures. The results obtained have proved that the proposed system performs better than ResNet-v2 (even with pretraining), and also better than other ViTs, which incorporate additional local information of patches with the LIFE module. We conclude that, if longer image datasets (or even from other modalities) about heart failures were available, overfitting would be reduced and robustness would be enhanced.
@@ Acknowledgements
The data that have been used are confidential. For further details, please contact with Enrique Lara-Pezzi (elara@cnic.es)).

### v2

@@ 5 Conclusions
Heart diseases are the main cause of human mortality in the world. Thus, the early identification of heart diseases is a task of great importance. In this work, a real-time heart failure prediction system for echocardiography video sequences has been presented to specifically address this challenging and demanding task, which has not been addressed in the related literature to the best of the authors’ knowledge. The two major contributions of this paper are the creation of a large annotated cardiac database from echocardiography sequences with heterogeneous acquisition features; and a deep neural network, based on a ViT, trained using a self-supervised learning approach which effectively improves the performance with scarce datasets. These two contributions, acting synergistically, successfully address the usual challenge of having a scarce number of samples in the medicine field, reducing the dependence on very large training datasets. More precisely, the elaboration of varied databases with high quality enough is very difficult, implying high costs and very hard specialized work from experts in heart diseases. In addition, the general trend in the related literature about heart disease recognition and heart failure prediction in echocardiography images barely explores self-supervised learning, even less Masked Autoencoders (MAE), to deal with this typical situation of scarce databases. This implies the necessity to adopt techniques specifically designed to address this problem. In the creation process of the database, the HODMD algorithm has been employed as a feature extractor and as a data augmentation technique. This algorithm has demonstrated to improve the heart failure time prediction performance, as also incorporating temporal information from the HODMD modes, useful to better characterize the different heart states and so to more accurately estimate the times of heart failures. The results obtained have proved that the proposed system performs better than ResNet-v2 (even with pretraining), and also better than other ViTs, which incorporate additional local information of patches with the LIFE module. We conclude that, if longer image datasets, and also from other modalities, about heart failures were available, overfitting would be reduced and robustness would be enhanced.
@@ Acknowledgements
The data that have been used are confidential. For further details, please contact with Enrique Lara-Pezzi (elara@cnic.es).